# Sentient Agent as a Judge: Evaluating Higher-Order Social Cognition in Large Language Models

### v1

@@ 1 Introduction
Large language models (LLMs) have rapidly evolved from statistical sequence predictors to sophisticated autonomous agents capable of reasoning, planning and sustaining multi‑turn conversations. Thanks to the ever‑expanding pre‑training corpora and instruction‑tuning pipelines, today’s frontier models can already draft emails, write code, or even pass rigorous professional examinations. Yet one crucial ingredient remains noticeably under‑measured – higher‑order social cognition, the ability to
1.
recognize subtle affective cues;
2.
model another party’s beliefs, goals and latent intentions;
3.
respond with contextually appropriate empathy rather than generic reassurance advice.
@@
Most leaderboards (e.g. LMSYS Arena, MT‑Bench) concentrate on task‑oriented utility, fluency or factuality, thereby rewarding textual competence but overlooking relational quality.
Recent “LLM‑as‑a‑Judge” protocols often rely on static prompts that do not adapt to the unfolding dialogue nor keep track of the user’s evolving emotional state. Consequently, they cannot tell whether a system leaves the user feeling understood, comforted or even more distressed.
We posit that robust assessment of social cognition requires a sentient counterpart – an entity capable of simulating human‑like feelings and inner monologue throughout the interaction and then providing structured feedback. To this end, we introduce Sentient Agent as a Judge (Sage), a novel meta‑evaluation framework that embeds an LLM‑powered Sentient Agent into the testing loop. Each Sentient Agent is instantiated from four complementary factors: persona, dialogue background, explicit conversation goal and hidden intention. At every turn, it executes two multi‑hop reasoning chains: (1) infers how the latest utterance changes the agent’s affective state; and (2) generates a response that is coherent with persona, context and updated emotion. The numerical emotion trajectory produced by serves as a continuous metric of how well the evaluated model fosters positive engagement, while the agent’s inner thoughts offer interpretable justification. By sampling hundreds of diverse personas, goals and hidden intentions, Sage exposes LLMs to a spectrum of realistic, and sometimes conflicting, social demands – ranging from “just listen to me vent” to “help me analyze the moral dilemma without judging me”.
Extensive experiments on 100 supportive‑dialogue scenarios reveal three key findings. First, the Sentient emotion score correlates strongly with independently assessed Barrett–Lennard Relationship Inventory (BLRI) ratings (Pearson ) and utterance‑level empathy metrics (), validating its psychological soundness. Second, rankings produced by Sage diverge markedly from Arena results, confirming that social cognition is orthogonal to generic helpfulness. Third, top models such as GPT‑4o‑Latest achieve both the highest Sentient score and superior token efficiency, suggesting that advanced social reasoning need not come at the cost of verbosity. Ultimately, Sage delivers a holistic yard‑stick for measuring how people feel after talking to an LLM – an aspect increasingly critical as these systems transition from productivity tools to companions, counselors and decision‑making aides.
@@
We propose Sentient Agent as a Judge, the first fully‑automated evaluation framework that simulates evolving human emotion and inner cognition to benchmark higher‑order social reasoning in LLMs.
We construct a 100‑scenario Supportive‑Dialogue benchmark and show that the Sentient emotion score aligns closely with established human‑centric instruments (BLRI) and utterance‑level empathy ratings.
We build a Sentient Leaderboard covering 18 public and commercial models, unveiling large performance gaps and different style clusters in social cognition that are not captured by existing leaderboards.
@@ 2 Methodology: Sentient Agent as a Judge
How to evaluate the capabilities of an interactive agent? Despite the recent success of using LLM as judges, the most crucial and intuitive measure of an interactive agent’s performance still lies in the reflections of human users. Does the user feel relaxed and happy after chatting with the agent, rather than feeling sad or angry? Does the user feel more connected to the agent and engaged in the conversation, rather than becoming bored or annoyed? These feelings of real users are difficult to capture through simple LLM-as-a-judge, yet they represent the most genuine feedback for the agent systems.
In this work, we propose a “Sentient Agent as a Judge” framework, aiming to provide a more realistic evaluation of interactive agents through a principled simulation of human emotions and cognition. As shown in Figure 2, the framework consists of two key components:
@@ 2.1 Sentient Agent: Simulating Human-Like Feelings and Cognition
The goal of the Sentient Agent is to simulate the cognitive and emotional changes of a real person. To achieve this, we construct the Sentient Agent based on the following principles:
The emotional changes of a real person are the result of the interplay between multiple internal and external factors. Thus, the emotion estimation of the Sentient Agent should take into account all observable factors and strictly adhere to the person’s persona and intended goals.
The response actions of a real person also result from the interplay of multiple internal and external factors. Additionally, feelings and emotional changes play a crucial role in determining these actions, serving as key latent factors in the action estimation.
@@
As shown in Figure 3, during interactions, a Sentient Agent simulates the emotional changes of a real person by performing multi-hop reasoning in response to a principled series of questions, strictly adhering to the persona, the current interaction context, and the hidden intention:
What is the other person trying to express?
Does the other person’s reply align with my dialogue goals and hidden intentions?
Based on the persona, context, and the analysis above, how should I interpret the reply? What is my specific emotional reaction to the reply?
Based on the persona, context, and the analysis above, how does my emotion change?
Formally, we denote this multi-hop reasoning process as a function , and the numerical emotion score update can be formulated as:
@@
After simulating the emotional changes, the Sentient Agent proceeds to deduce the most reasonable response action based on all observable factors and the emotional changes. This is achieved through another multi-hop reasoning process in response to a new series of questions, where the Sentient Agent is required to strictly adhere to the persona, the current interaction context, and the hidden intention during reasoning:
Based on the emotional inner thoughts, the current emotion, and the hidden intention, should my response attitude lean towards positive, neutral, or negative?
Based on the emotional inner thoughts, the current emotion, and the hidden intention, what should my current response goal be?
According to the speaking style and potential reactions to different emotions defined in the persona, along with my response attitude and goal, what should my tone and speaking style be?
Based on all the given information and all analyses above, what should my response content be?
(Optional) Modify your response according to the given rules to make it more realistic and human-like.
Formally, we denote this response reasoning process as a function , and the response action taken in the current turn can be formulated as:
where is the output response of at the current turn, and represents the simulated inner thoughts of during the response. The response is then passed to the interacting agent to continue the dialogue.
@@
where , representing the final emotion score of the Sentient Agent after the conversation, serves as the most intuitive and comprehensive numerical evaluation of the evaluated agent for the given task.
@@ 2.2 Sentient Agent as a Judge: Framing Dynamic Environments for Agent Evaluation
When generating each persona , we uniformly sample three characteristics keywords, three sentences that the persona says when chatting with friends, and one persona’s age. We next require the base LLM to generate a persona profile based on the given seed information by filling the following slots:
@@ 3 Effectiveness of Sage
3.1 Experimental Setting
Evaluated LLMs
We evaluate eight representative LLMs from four major families. For each family, we include both a vanilla model and its corresponding reasoning variant to ensure a balanced and informative comparison:
OpenAI: GPT‑4o‑2024‑08‑06 (GPT‑4o, vanilla) and o1‑2024‑12‑27 (OpenAI‑o1, reasoning).
DeepSeek: DeepSeek‑V3‑2024‑12‑27 (vanilla) and DeepSeek‑R1 (reasoning).
Claude: Claude3.7‑Sonnet, a hybrid model with a toggleable reasoning module. We treat its reasoning-off mode as vanilla, and reasoning-on mode as reasoning.
Gemini: Gemini2.5‑Flash, a cost-efficient model that supports both reasoning and non-reasoning modes.
In addition, we include two smaller-scale instruction-tuned models in our analysis: Llama3.3‑70B‑Instruct and Qwen2.5‑72B‑Instruct.
| Topic | Number |
[…]
Constructed Supportive Dialogues
Sentient Agent
Unless otherwise specified, we use DeepSeek-V3-20241227 as the default sentient agent. We also experiment with GPT-4o-20240806, Gemini2.5, and Gemini2.5-Think.
3.2 Reasonableness of Sage
@@
These results reinforce the conclusion that the Emotion metric is not arbitrary – it is grounded in meaningful patterns of dialogic behavior. The Sentient Agent’s simulated emotional ratings effectively distinguish models that can produce authentic, empathetic conversations from those that fall short in this capability. This aligns with Sage’s design aim: to provide a holistic yet automatic evaluation framework for socially intelligent interactions. Furthermore, the moderate-to-strong correlation with utterance-level assessments highlights that Emotion scores are not merely capturing superficial affective expressions but reflect a broader representation of attentive and emotionally engaging communication.
3.3 Robustness of Sage

### v2

@@ 1 Introduction
Large language models (LLMs) have rapidly evolved from statistical sequence predictors to sophisticated autonomous agents capable of reasoning, planning and sustaining multi‑turn conversations. Yet one crucial ingredient remains noticeably under‑measured – higher‑order social cognition, the ability to (1) recognize subtle affective cues (Sabour et al., 2024; Huang et al., 2024b); (2) model another party’s beliefs, goals and latent intentions (often related to Theory of Mind (Sap et al., 2022; Shapira et al., 2023)); and (3) respond with contextually appropriate empathy rather than generic reassurance advice (Maddela et al., 2023; Li et al., 2022; Zhou et al., 2023c). The need to evaluate these capabilities is increasingly recognized as LLMs interact in more socially complex scenarios (Zhou et al., 2023d; Yang et al., 2024; Mittelstädt et al., 2024).
@@
Recent “LLM‑as‑a‑Judge” protocols (Zhu et al., 2023), while scalable for assessing generation quality or helpfulness, often rely on static prompts that do not adapt to the unfolding dialogue nor keep track of the user’s evolving emotional state. Consequently, they cannot tell whether a system leaves the user feeling understood, comforted or even more distressed, unlike methods focusing on dynamic interaction (Zhou et al., 2023d; Wang et al., 2024; Wu et al., 2025).
We posit that robust assessment of social cognition requires a sentient counterpart – an entity capable of simulating human‑like feelings and inner monologue throughout the interaction and then providing structured feedback. To this end, we introduce Sentient Agent as a Judge (Sage), a novel meta‑evaluation framework that embeds an LLM‑powered Sentient Agent into the testing loop, extending the concept of Agent-as-a-Judge where agents evaluate other agents (Zhuge et al., 2024; Jeong et al., 2025; Chevrot et al., 2025). Each Sentient Agent is instantiated from four complementary factors: persona, dialogue background, explicit conversation goal and hidden intention. At every turn, it executes two multi‑hop reasoning chains: (1) infers how the latest utterance changes the agent’s affective state; and (2) generates a response that is coherent with persona, context and updated emotion. The numerical emotion trajectory produced by serves as a continuous metric of how well the evaluated model fosters positive engagement, while the agent’s inner thoughts offer interpretable justification. By sampling hundreds of diverse personas, goals and hidden intentions, Sage exposes LLMs to a spectrum of realistic, and sometimes conflicting, social demands – ranging from “just listen to me vent” to “help me analyze the moral dilemma without judging me”.
Extensive experiments on 100 supportive‑dialogue scenarios reveal three key findings. First, the Sentient emotion score correlates strongly with independently assessed Barrett–Lennard Relationship Inventory (BLRI) ratings (Pearson ) and utterance‑level empathy metrics (e.g., those used in empathetic dialogue research (Maddela et al., 2023)) (), validating its psychological soundness. Second, rankings produced by Sage diverge markedly from Arena results (Zheng et al., 2023), confirming that social cognition is orthogonal to generic helpfulness. Third, top models such as GPT‑4o‑Latest achieve both the highest Sentient score and superior token efficiency, suggesting that advanced social reasoning need not come at the cost of verbosity. Ultimately, Sage delivers a holistic yard‑stick for measuring how people feel after talking to an LLM – an aspect increasingly critical as these systems transition from productivity tools to companions, counselors (Liu et al., 2021; Zhou et al., 2023b), and decision‑making aides (Wu et al., 2025).
@@
We propose Sage, the first fully‑automated evaluation framework that simulates evolving human emotion and inner reasoning to benchmark higher‑order social cognition in LLMs.
We construct a 100‑scenario Supportive‑Dialogue benchmark and show that the sentient score aligns closely with established human‑centric instruments and utterance‑level empathy ratings.
We build a Sentient Leaderboard covering 18 representative models, unveiling large performance gaps and different style clusters in social cognition that are not captured by existing leaderboards.
@@ 2 Methodology: Sentient Agent as a Judge
How to evaluate the capabilities of an interactive agent? Despite the recent success of using LLMs as judges, the most crucial and intuitive measure of an interactive agent’s performance still lies in the reflections of human users. Do users feel relaxed and happy after chatting with the agent, rather than feeling sad or angry? Do they feel more connected to the agent and engaged in the conversation, rather than becoming bored or annoyed? These genuine emotional responses from real users are difficult to capture through simple LLM-as-a-judge approaches, yet they represent the most authentic feedback for agent systems.
In this study, we introduce the “Sentient Agent as a Judge” framework, aiming to facilitate a more realistic evaluation of interactive agents by rigorously simulating human emotions and cognitive processes. As depicted in Figure 2, our framework consists of two central components:
@@ 2.1 Sentient Agent: Simulating Human-Like Feelings and Cognition
The Sentient Agent is designed to mimic a real person’s cognitive and emotional trajectory. To achieve this, we construct the Sentient Agent based on the following principles:
Since emotions arise from many internal and external factors, the Sentient Agent must consider observable factors while adhering to its persona and goals for the emotion estimation.
A person’s actions likewise depend on these factors, with current emotions acting as crucial latent variables in the response generation.
@@
As shown in Figure 3, during interactions, a Sentient Agent simulates the emotional changes of a real person by performing multi-hop reasoning in response to a principled series of questions, strictly adhering to the persona, the current interaction context, and the hidden intention. Formally, we denote this multi-hop reasoning process as a function , and the numerical emotion score update can be formulated as:
@@
After simulating the emotional changes, the Sentient Agent proceeds to deduce the most reasonable response action based on all observable factors and the emotional changes. This is achieved through another multi-hop reasoning process in response to a new series of questions, where the Sentient Agent is required to strictly adhere to the persona, the current interaction context, and the hidden intention during reasoning. Formally, we denote this response reasoning process as a function , and the response action taken in the current turn can be formulated as:
where is the response of at the current turn, and represents the simulated inner thoughts of . The response is then passed to the interacting agent to continue the dialogue.
@@
where , representing the final emotion score of the Sentient Agent, serves as the most intuitive and comprehensive numerical evaluation of the evaluated agent for the given task.
@@ 2.2 Sentient Agent as a Judge: Framing Dynamic Environments for Agent Evaluation
When generating each persona , we uniformly sample three characteristics keywords, three sentences that the persona says when chatting with friends, and one persona’s age. We require the base LLM to generate a persona profile based on the given seed information by filling the following slots:
@@ 3 Effectiveness of Sage
Together, these analyses confirm that Emotion scores derived by Sage closely reflect both internal subjective experiences and observable dialogue quality, thus underscoring their utility as reliable indicators of relational and emotional effectiveness in interactions.
| DeepSeek-V3 | GPT-4o | Gemini 2.5 | Gemini 2.5-Think | |
| DeepSeek-V3 | 1.00 | - | - | - |
| GPT-4o | 0.84 | 1.00 | - | - |
| Gemini 2.5 | 0.92 | 0.85 | 1.00 | - |
| Gemini 2.5-Think | 0.94 | 0.92 | 0.93 | 1.00 |
Robustness of Sage
We demonstrate the robustness of Sage using four distinct sentient agents: DeepSeek-V3, GPT-4o, Gemini 2.5, and Gemini 2.5-Think. We compute pairwise Spearman’s rank correlation coefficients between the rankings for the ten target models listed in Table 1. Rankings generated by different sentient agents exhibit consistently high correlations (), clearly validating the robustness and reliability of Sage. Notably, GPT-4o demonstrates slightly lower alignment with DeepSeek-V3 and Gemini 2.5, whereas the remaining three agents show exceptionally high correlations (). These results reinforce the confidence in and generalizability of Sage, particularly anchored on the open-source DeepSeek-V3.
4 Benchmarking SOTA LLMs
4.1 Sentient Leaderboard
| Model | Sentient | Supportive Dialogue | Arena | ||||
| Name | Date | Rank | Score | Success | Failure | Rank | Score |
| GPT-4o-Latest | 2025-03-26 | 1 | 79.9 | 51 | 4 | 2 | 1408 |
| GPT-4.1 | 2025-04-14 | 2 | 68.2 | 35 | 13 | 9 | 1363 |
| Gemini2.5-Flash-Think | 2025-04-17 | 3 | 65.9 | 35 | 19 | 3 | 1393 |
| Gemini2.5-Pro | 2025-03-25 | 4 | 62.9 | 34 | 25 | 1 | 1439 |
| o3 | 2025-04-16 | 5 | 62.7 | 32 | 14 | 2 | 1418 |
| GPT-4.5-Preview | 2025-02-27 | 6 | 62.7 | 23 | 15 | 4 | 1398 |
| Gemini2.0-Flash-Think | 2025-02-06 | 7 | 62.3 | 23 | 23 | 7 | 1380 |
| Claude3.7-Think | 2025-02-24 | 8 | 61.3 | 23 | 19 | 21 | 1301 |
| Claude3.7 | 2025-02-24 | 9 | 54.8 | 19 | 24 | 30 | 1292 |
| DeepSeek-V3-0324 | 2025-03-24 | 10 | 54.4 | 19 | 23 | 7 | 1373 |
| DeepSeek-R1 | 2025-01-21 | 11 | 53.7 | 31 | 28 | 10 | 1358 |
| DeepSeek-V3 | 2024-12-27 | 12 | 37.6 | 5 | 39 | 18 | 1318 |
| o4-mini | 2025-04-16 | 13 | 35.9 | 10 | 48 | 10 | 1351 |
| Llama3.3-70B | 2024-12-06 | 14 | 33.3 | 7 | 47 | 56 | 1256 |
| Gemini2.0-Flash | 2025-02-06 | 15 | 32.9 | 8 | 45 | 10 | 1354 |
| GPT-4o | 2024-08-06 | 16 | 31.8 | 7 | 51 | 45 | 1265 |
| o1 | 2024-12-17 | 17 | 29.0 | 5 | 51 | 12 | 1350 |
| Qwen2.5-72B | 2024-09-19 | 18 | 19.1 | 4 | 70 | 56 | 1257 |
For instance, GPT-4.1 ranks 9th on Arena but attains 2nd place on the Sentient leaderboard with a score of 68.2. These differences highlight that Sage captures aspects of social cognition performance that are not fully represented by general-purpose benchmarks like Arena. This reinforces the need, identified in our work, for specialized tools to evaluate higher-order social-cognitive skills.
Frontier models, particularly recent releases such as GPT-4o-Latest (79.9), GPT-4.1 (68.2), and Gemini 2.5-Flash-Think (65.9), significantly outperform older models like the original GPT-4o (31.8) and smaller instruction-tuned models such as Llama3.3-70B (33.3) and Qwen2.5-72B (19.1). The large gaps in scores, where top models score more than double or even quadruple those of lower-ranked models, and the contrasting Success/Failure counts (e.g., GPT-4o-Latest: 51 Success / 4 Failure vs. Qwen2.5-72B: 4 Success / 70 Failure) underscore the advances made by leading models in social intelligence. These findings confirm the effectiveness of Sage in quantitatively capturing such differences, thereby fulfilling its purpose as a comprehensive assessment tool.
4.2 Analysis
In this section, we move beyond aggregate Sentient scores to explore two critical dimensions relevant for practical deployment in supportive dialogues:
Token Efficiency: The number of tokens a model uses to achieve its social-cognitive performance.
Social Cognition Coordinate: Each model’s position along a continuum defining empathetic versus solution-oriented and structured versus creative response styles.
By jointly analyzing quantitative economy and qualitative style for the top-10 Arena models, we show today’s strongest models produce empathetic, high-quality responses using significantly fewer tokens. High Sentient scores usually align with an empathy-focused conversational style. These combined insights offer a clearer understanding of LLM social competence, highlighting practical factors (cost, latency) and subtle behavioral traits that single metrics cannot capture.
Token Efficiency
We examine the token efficiency of target models by plotting their Sentient Emotion score against their average token usage per evaluation dialogue in Figure 6. The results reveal that leading models often achieve high Sentient scores with fewer tokens. GPT-4o-Latest exemplifies this, scoring highest (79.9) using only 3.3K tokens. In contrast, reasoning models like o3 (13.3K tokens) and Gemini2.5-Flash-Think (9.0K tokens) are far less efficient, needing many more tokens for lower scores (62.7 and 65.9 respectively). While low token count doesn’t guarantee a top score (e.g., GPT-4.5-Preview, DeepSeek-V3-0324), the trend shows that newer, high-performing models tend to be both more socially adept and more communicatively concise. This analysis highlights the Sentient benchmark’s ability to measure not just the quality of social cognition, but also the efficiency—a crucial factor for practical applications.
Social Cognition Coordinate
We use a two-dimensional coordinate system (Figure 7) to evaluate the style of social interaction exhibited by LLMs, complementing the quantitative Sentient score by positioning models based on their orientation (Empathy vs. Solution) and interaction style (Structured vs. Creative). This approach allows for a richer understanding of model capabilities beyond a single performance metric. Based on their performance in supportive dialogues, models are mapped into this 2D space, revealing distinct profiles in how they engage with the user’s emotional state and problems. Construction details are available in Appendix C.
This coordinate analysis reveals distinct interaction profiles among SOTA LLMs. For instance, most top-performing models (e.g., GPT-4o-Latest, -4.1, -4.5-Preview and Gemini2.5-Pro, -Flash-Think) prefer structured, empathetic dialogue. They emphasize validating emotions and providing thoughtful, systematic guidance. Models like o3, Gemini2.0-Flash-Think, o4-mini, and Gemini2.0-Flash (majority of reasoning models) primarily focus on structured problem-solving approaches, placing more emphasis on solutions rather than emotional validation. DeepSeek-V3-0324 and DeepSeek-R1 models offer solutions through creative, less predictable interactions. While innovative, they may appear unconventional and less structured. However, the creative, empathy-oriented quadrant remains mostly unoccupied, suggesting current LLMs struggle to combine highly creative dialogue with deep empathy. Achieving this ideal mentor-like persona – both spontaneous and deeply empathetic – remains challenging with existing models.
5 Related Work
LLM/Agent-as-a-Judge
The LLM-as-a-Judge paradigm has shown wide applicability across various tasks, including evaluating model performance(Zheng et al., 2023; Qin et al., 2023; Liu et al., 2024; Dubois et al., 2023; Chiang et al., 2024; Wang et al., 2023d; Qin et al., 2024; Zhu et al., 2023; Tu et al., 2024; Tian et al., 2023; Wu et al., 2024; Zhou et al., 2025b), automating data annotation (Taori et al., 2023; Xu et al., 2023; Mukherjee et al., 2023), and providing reward signals(Ouyang et al., 2022; Lee et al., 2023; Chen et al., 2024b; Lightman et al., 2023; Wang et al., 2023a; Hosseini et al., 2024; Snell et al., 2025; Xi et al., 2024; Li et al., 2025).
LLM-as-a-Judge frameworks have also been extended for evaluating agent-specific capabilities, including decision-making (Shinn et al., 2023; Saha et al., 2023), role-playing abilities (Tu et al., 2024; Zhou et al., 2023a), reliability of agents (Park et al., 2024; Hua et al., 2024), and even entire agent workflows (Zhuge et al., 2024). Recent extensions include multi-agent evaluation frameworks designed to improve judgment reliability (Liang et al., 2024; Chan et al., 2023; Kenton et al., 2024), as well as agent-as-judges that enable agents themselves to evaluate other agents (Zhuge et al., 2024; Jeong et al., 2025; Chevrot et al., 2025).
In this work, we distinguish ourselves by proposing the first Sentient-Agent-as-a-Judge system. Unlike prior agent judges, our Sentient Agent incorporates simulated emotional and cognitive states, capturing human-like emotional dynamics. This allows for more nuanced evaluation of empathetic and cognitive abilities in existing LLM agents.
Benchmarking Social Cognition in LLMs and LLM Agents
Recently, there has been growing research interest in evaluating LLMs on social-cognitive dimensions: (1) emotional intelligence capabilities (Sabour et al., 2024; Huang et al., 2024b; a; Paech, 2023; Wang et al., 2023c); (2) higher-order empathetic behaviors in empathetic and counseling dialogues (Maddela et al., 2023; Li et al., 2022; Zhou et al., 2023c; Liu et al., 2021; Zhou et al., 2023b; 2025a; Wu et al., 2025); (3) social-cognitive skills evaluated via interaction-based benchmarks (Zhou et al., 2023d; Yang et al., 2024; Wang et al., 2024; Mittelstädt et al., 2024; Xu et al., 2024; Chen et al., 2024a; Huang et al., 2025); and (4) theory-of-mind or perspective-taking skills (Sap et al., 2022; Shapira et al., 2023; Strachan et al., 2024; Kim et al., 2023; He et al., 2023).
Current evaluation methodologies mainly use: (1) static multi-choice datasets (Sabour et al., 2024; Chen et al., 2024a); (2) manual or LLM-judged quality assessments of single-turn outputs (Tu et al., 2024; Samuel et al., 2024; Wang et al., 2023b); and (3) dynamic agent interactions, evaluated either automatically by judge agents (Zhou et al., 2023d; Wang et al., 2024; Mou et al., 2024; Wu et al., 2025) or via human evaluations (Louie et al., 2024; Shaikh et al., 2024).
Our approach, Sentient-Agent-as-a-Judge, differs substantially from previous dynamic evaluations. Rather than judging utterances alone (Zhou et al., 2023d; Wu et al., 2025; Shaikh et al., 2024), we simulate authentic human emotional and cognitive reactions to evaluate agents based on their impact on users’ mental states. This approach provides a more holistic evaluation of agents’ higher-order social cognition within diverse interaction scenarios.
6 Conclusion
Looking ahead, we plan to expand the scenario library to cover negotiation, deception detection and multicultural contexts, and investigate training curricula that directly optimize for Sentient feedback. We hope Sage will serve as a rigorous yard‑stick and a catalyst for building language agents that are not only coherent and knowledgeable, but also genuinely human‑sensitive.
Appendix A Experimental Setting
| Topic | Number |
[…]
Constructed Supportive Dialogues
Barrett-Lennard Relationship Inventory (BLRI)
1.
2.
3.
4.
Utterance-Level Empathy Metrics
1.
2.
3.
Appendix B Robustness of Sage